### v2

@@ A.2 Proofs of Section 8
$-\int^{t}_{0}$ $\sum^{n}_{i=1}\int_{\mathbb{R}^{n}}\left(\sum^{n}_{j=1}\frac{1}{2}\frac{\partial a^{\alpha}_{ij}}{\partial y_{j}}(s,y)-b_{i}(s,y)\right)\frac{\partial}{\partial y_{i}}\left(\frac{\partial p_{k}}{\partial s}\right)(s,y)\ p_{k}(s,y)\;\mathrm{d}y\;\mathrm{d}s=$
$=\int^{t}_{0}\sum^{n}_{i=1}\int_{\mathbb{R}^{n}}\left(\sum^{n}_{j=1}\frac{1}{2}\frac{\partial^{2}a^{\alpha}_{ij}}{\partial y_{i}\partial y_{j}}(s,y)-\frac{\partial b_{i}}{\partial y_{i}}(s,y)\right)\frac{\partial p_{k}}{\partial s}(s,y)p_{k}(s,y)\;\mathrm{d}y\;\mathrm{d}s$
$\quad+\int^{t}_{0}\sum^{n}_{i=1}\int_{\mathbb{R}^{n}}\left(\sum^{n}_{j=1}\frac{1}{2}\frac{\partial a^{\alpha}_{ij}}{\partial y_{j}}(s,y)-b_{i}(s,y)\right)\frac{\partial p_{k}}{\partial y_{i}}(s,y)\frac{\partial p_{k}}{\partial s}(s,y)\;\mathrm{d}y\;\mathrm{d}s$
$\leq C\big{(}\alpha,\|(a,b)\|_{\mathcal{H}_{m}}\big{)}\int^{t}_{0}\int_{\mathbb{R}^{n}}\left(|p_{k}(s,y)|+\sum^{n}_{i=1}\left|\frac{\partial p_{k}}{\partial y_{i}}(s,y)\right|\right)\left|\frac{\partial p_{k}}{\partial s}(s,y)\right|\;\mathrm{d}y\;\mathrm{d}s$
$\leq\frac{1}{2}\int^{t}_{0}\left\|\frac{\partial p_{k}}{\partial s}(s,\cdot)\right\|^{2}_{L^{2}}\;\mathrm{d}s+C\big{(}\alpha,\|(a,b)\|_{\mathcal{H}_{m}}\big{)}\left(\|\bar{p}\|^{2}_{H^{1}}+\int^{t}_{0}\|f(s,\cdot)\|^{2}_{L^{2}}\;\mathrm{d}s\right),$
@@
(38) $\int^{t}_{0}\left\|\frac{\partial p_{k}}{\partial s}(s,\cdot)\right\|^{2}_{L^{2}}\;\mathrm{d}s\leq C\big{(}\alpha,\|(a,b)\|_{\mathcal{H}^{+}_{m}}\big{)}\left(\|\bar{p}\|^{2}_{H^{1}}+\int^{t}_{0}\|f(s,\cdot)\|^{2}_{L^{2}}\;\mathrm{d}s\right).$
@@
(40) $\int_{\mathbb{R}^{n}}\sum^{n}_{i,j=1}\frac{1}{2}a^{\alpha}_{ij}(t,y)\frac{\partial p}{\partial y_{j}}(t,y)\frac{\partial\varphi}{\partial y_{i}}(y)\;\mathrm{d}y\hskip 86.11084pt\textnormal{for}\quad\varphi\in H^{1}(\mathbb{R}^{n},\mathbb{R})$
$+\sum^{n}_{i=1}\int_{\mathbb{R}^{n}}\left(\sum^{n}_{j=1}\frac{1}{2}\frac{\partial a^{\alpha}_{ij}}{\partial y_{j}}(t,y)-b_{i}(t,y)\right)\frac{\partial\varphi}{\partial y_{i}}(y)p(t,y)\;\mathrm{d}y=\int_{\mathbb{R}^{n}}g(t,y)\varphi(y)\;\mathrm{d}y,$
@@
(42) $\|p(t,\cdot)\|^{2}_{H^{2}}\leq C\big{(}\alpha,\|(a,b)\|_{\mathcal{H}_{m}}\big{)}\Big{(}\|p(t,\cdot)\|^{2}_{L^{2}}+\|g(t,\cdot)\|^{2}_{L^{2}}\Big{)}.$
@@
(44) $\leq C\big{(}\alpha,\|(a,b)\|_{\mathcal{H}_{m}}\big{)}\left(\|\bar{p}\|^{2}_{H^{1}}+\int^{t}_{0}\|f(s,\cdot)\|^{2}_{L^{2}}\;\mathrm{d}s\right),\quad t\in[0,T].$
@@
$A^{0,\dot{a},\dot{b}}(t;p(t,\cdot),\varphi)=-\int_{\mathbb{R}^{n}}(\mathcal{L}^{0,\dot{a},\dot{b}}_{t})^{*}p(t,y)\varphi(y)\;\mathrm{d}y,$
for every $\varphi\in H^{1}(\mathbb{R}^{n},\mathbb{R})$ and almost every $t\in[0,T]$. Therefore, since $(\mathcal{L}^{0,\dot{a},\dot{b}}_{t})^{*}p\in L^{2}([0,T]\times\mathbb{R}^{n},\mathbb{R})$ and $(\mathcal{L}^{a,b}_{0})^{*}\bar{p}\in H^{1}(\mathbb{R}^{n},\mathbb{R})$, problem (45) fits the setting of Theorem 8.9, thus from our previous computations there exists a unique solution $q\in L^{2}(0,T;H^{2}(\mathbb{R}^{n},\mathbb{R}))$ to (45), which additionally satisfies $\frac{\partial q}{\partial t}\in L^{2}([0,T]\times\mathbb{R}^{n},\mathbb{R})$.
@@
$\quad-\sum^{n}_{i,j=1}\int_{\mathbb{R}^{n}}\frac{1}{2}\left(\int^{t}_{0}\left(a^{\alpha}_{ij}(s,y)\frac{\partial}{\partial s}\left(\frac{\partial w}{\partial y_{j}}\right)(s,y)+\dot{a}^{\alpha}_{ij}(s,y)\frac{\partial w}{\partial y_{j}}(s,y)\right)\mathrm{d}s\right)\frac{\partial\varphi}{\partial y_{i}}(y)\;\mathrm{d}y$
$\quad-\sum^{n}_{j=1}\int_{\mathbb{R}^{n}}\Bigg{(}\int^{t}_{0}\Bigg{(}\Bigg{(}\sum^{n}_{j=1}\frac{\partial\dot{a}^{\alpha}_{ij}}{\partial y_{j}}(s,y)-\dot{b}_{i}(s,y)\Bigg{)}w(s,y)$
$\hskip 123.78433pt+\Bigg{(}\sum^{n}_{j=1}\frac{\partial a^{\alpha}_{ij}}{\partial y_{j}}(s,y)-b_{i}(s,y)\Bigg{)}\frac{\partial w}{\partial s}(s,y)\Bigg{)}\mathrm{d}s\Bigg{)}\frac{\partial\varphi}{\partial y_{i}}(y)\;\mathrm{d}y$
@@
$t\mapsto\int^{t}_{0}(\mathcal{L}^{0,\dot{a},\dot{b}}_{s})^{*}(z_{1}-z_{2})(s,\cdot)\;\mathrm{d}s,$
@@
(48) $\hskip 21.52771pt\leq C\big{(}\alpha,\|(a,b)\|_{\mathcal{H}_{m}}\big{)}\int^{t}_{0}\left(\int^{s}_{0}\|(z_{1}-z_{2})(r,\cdot)\|^{2}_{H^{2}}\;\mathrm{d}r\right)\;\mathrm{d}s,\quad t\in[0,T],$
@@
$\int^{t}_{0}\|(z_{1}-z_{2})(s,\cdot)\|^{2}_{H^{2}}\;\mathrm{d}s$ $\leq C\big{(}\alpha,\|(a,b)\|_{\mathcal{H}_{m}}\big{)}\int^{t}_{0}\|h(s,\cdot)\|^{2}_{L^{2}}\;\mathrm{d}s$
$\leq C\big{(}\alpha,\|(a,b)\|_{\mathcal{H}_{m}}\big{)}\int^{t}_{0}\left(\int^{s}_{0}\|(z_{1}-z_{2})(r,\cdot)\|^{2}_{H^{2}}\;\mathrm{d}r\right)\;\mathrm{d}s,$

### v3

@@ A.2 Proofs of Section 8
$-\int^{t}_{0}\sum^{n}_{i=1}\int_{\mathbb{R}^{n}}\left(\sum^{n}_{j=1}\frac{1}{2}\frac{\partial a_{ij}}{\partial y_{j}}(s,y)-b_{i}(s,y)\right)\frac{\partial}{\partial y_{i}}\left(\frac{\partial p_{k}}{\partial s}\right)(s,y)\ p_{k}(s,y)\;\mathrm{d}y\;\mathrm{d}s=$
$=\int^{t}_{0}\sum^{n}_{i=1}\int_{\mathbb{R}^{n}}\left(\sum^{n}_{j=1}\frac{1}{2}\frac{\partial^{2}a_{ij}}{\partial y_{i}\partial y_{j}}(s,y)-\frac{\partial b_{i}}{\partial y_{i}}(s,y)\right)\frac{\partial p_{k}}{\partial s}(s,y)p_{k}(s,y)\;\mathrm{d}y\;\mathrm{d}s$
$\quad+\int^{t}_{0}\sum^{n}_{i=1}\int_{\mathbb{R}^{n}}\left(\sum^{n}_{j=1}\frac{1}{2}\frac{\partial a_{ij}}{\partial y_{j}}(s,y)-b_{i}(s,y)\right)\frac{\partial p_{k}}{\partial y_{i}}(s,y)\frac{\partial p_{k}}{\partial s}(s,y)\;\mathrm{d}y\;\mathrm{d}s$
$\leq C\big{(}\alpha,\|({\color[rgb]{0,0,0}a-\alpha I},b)\|_{\mathcal{H}_{m}}\big{)}\int^{t}_{0}\int_{\mathbb{R}^{n}}\left(|p_{k}(s,y)|+\sum^{n}_{i=1}\left|\frac{\partial p_{k}}{\partial y_{i}}(s,y)\right|\right)\left|\frac{\partial p_{k}}{\partial s}(s,y)\right|\;\mathrm{d}y\;\mathrm{d}s$
$\leq\frac{1}{2}\int^{t}_{0}\left\|\frac{\partial p_{k}}{\partial s}(s,\cdot)\right\|^{2}_{L^{2}}\;\mathrm{d}s+C\big{(}\alpha,\|({\color[rgb]{0,0,0}a-\alpha I},b)\|_{\mathcal{H}_{m}}\big{)}\left(\|\bar{p}\|^{2}_{H^{1}}+\int^{t}_{0}\|f(s,\cdot)\|^{2}_{L^{2}}\;\mathrm{d}s\right),$
@@
(40) $\int^{t}_{0}\left\|\frac{\partial p_{k}}{\partial s}(s,\cdot)\right\|^{2}_{L^{2}}\;\mathrm{d}s\leq C\big{(}\alpha,\|({\color[rgb]{0,0,0}a-\alpha I},b)\|_{\mathcal{H}_{m}}\big{)}\left(\|\bar{p}\|^{2}_{H^{1}}+\int^{t}_{0}\|f(s,\cdot)\|^{2}_{L^{2}}\;\mathrm{d}s\right).$
@@
(42) $\int_{\mathbb{R}^{n}}\sum^{n}_{i,j=1}\frac{1}{2}a_{ij}(t,y)\frac{\partial p}{\partial y_{j}}(t,y)\frac{\partial\varphi}{\partial y_{i}}(y)\;\mathrm{d}y\hskip 86.11084pt\textnormal{for}\quad\varphi\in H^{1}(\mathbb{R}^{n},\mathbb{R})$
$+\sum^{n}_{i=1}\int_{\mathbb{R}^{n}}\left(\sum^{n}_{j=1}\frac{1}{2}\frac{\partial a_{ij}}{\partial y_{j}}(t,y)-b_{i}(t,y)\right)\frac{\partial\varphi}{\partial y_{i}}(y)p(t,y)\;\mathrm{d}y=\int_{\mathbb{R}^{n}}g(t,y)\varphi(y)\;\mathrm{d}y,$
@@
(44) $\|p(t,\cdot)\|^{2}_{H^{2}}\leq C\big{(}\alpha,\|({\color[rgb]{0,0,0}a-\alpha I},b)\|_{\mathcal{H}_{m}}\big{)}\Big{(}\|p(t,\cdot)\|^{2}_{L^{2}}+\|g(t,\cdot)\|^{2}_{L^{2}}\Big{)}.$
@@
(46) $\leq C\big{(}\alpha,\|({\color[rgb]{0,0,0}a-\alpha I},b)\|_{\mathcal{H}_{m}}\big{)}\left(\|\bar{p}\|^{2}_{H^{1}}+\int^{t}_{0}\|f(s,\cdot)\|^{2}_{L^{2}}\;\mathrm{d}s\right),\quad t\in[0,T].$
@@
$A^{0,\dot{a},\dot{b}}(t;p(t,\cdot),\varphi)=-\int_{\mathbb{R}^{n}}(\mathcal{L}^{\dot{a},\dot{b}}_{t})^{*}p(t,y)\varphi(y)\;\mathrm{d}y,$
for every $\varphi\in H^{1}(\mathbb{R}^{n},\mathbb{R})$ and almost every $t\in[0,T]$. Therefore, since $(\mathcal{L}^{\dot{a},\dot{b}}_{t})^{*}p\in L^{2}([0,T]\times\mathbb{R}^{n},\mathbb{R})$ and $(\mathcal{L}^{a,b}_{0})^{*}\bar{p}\in H^{1}(\mathbb{R}^{n},\mathbb{R})$, problem (47) fits the setting of Theorem 8.9, thus from our previous computations there exists a unique solution $q\in L^{2}(0,T;H^{2}(\mathbb{R}^{n},\mathbb{R}))$ to (47), which additionally satisfies $\frac{\partial q}{\partial t}\in L^{2}([0,T]\times\mathbb{R}^{n},\mathbb{R})$.
@@
$\quad-\sum^{n}_{i,j=1}\int_{\mathbb{R}^{n}}\frac{1}{2}\left(\int^{t}_{0}\left(a_{ij}(s,y)\frac{\partial}{\partial s}\left(\frac{\partial w}{\partial y_{j}}\right)(s,y)+\dot{a}_{ij}(s,y)\frac{\partial w}{\partial y_{j}}(s,y)\right)\mathrm{d}s\right)\frac{\partial\varphi}{\partial y_{i}}(y)\;\mathrm{d}y$
$\quad-\sum^{n}_{j=1}\int_{\mathbb{R}^{n}}\Bigg{(}\int^{t}_{0}\Bigg{(}\Bigg{(}\sum^{n}_{j=1}\frac{\partial\dot{a}_{ij}}{\partial y_{j}}(s,y)-\dot{b}_{i}(s,y)\Bigg{)}w(s,y)$
$\hskip 123.78433pt+\Bigg{(}\sum^{n}_{j=1}\frac{\partial a_{ij}}{\partial y_{j}}(s,y)-b_{i}(s,y)\Bigg{)}\frac{\partial w}{\partial s}(s,y)\Bigg{)}\mathrm{d}s\Bigg{)}\frac{\partial\varphi}{\partial y_{i}}(y)\;\mathrm{d}y$
@@
$t\mapsto\int^{t}_{0}(\mathcal{L}^{\dot{a},\dot{b}}_{s})^{*}(z_{1}-z_{2})(s,\cdot)\;\mathrm{d}s,$
@@
(50) $\hskip 21.52771pt\leq C\big{(}\alpha,\|({\color[rgb]{0,0,0}a-\alpha I},b)\|_{\mathcal{H}_{m}}\big{)}\int^{t}_{0}\left(\int^{s}_{0}\|(z_{1}-z_{2})(r,\cdot)\|^{2}_{H^{2}}\;\mathrm{d}r\right)\;\mathrm{d}s,\quad t\in[0,T],$
@@
$\int^{t}_{0}\|(z_{1}-z_{2})(s,\cdot)\|^{2}_{H^{2}}\;$ $\mathrm{d}s\leq C\big{(}\alpha,\|({\color[rgb]{0,0,0}a-\alpha I},b)\|_{\mathcal{H}_{m}}\big{)}\int^{t}_{0}\|h(s,\cdot)\|^{2}_{L^{2}}\;\mathrm{d}s$
$\leq C\big{(}\alpha,\|({\color[rgb]{0,0,0}a-\alpha I},b)\|_{\mathcal{H}_{m}}\big{)}\int^{t}_{0}\left(\int^{s}_{0}\|(z_{1}-z_{2})(r,\cdot)\|^{2}_{H^{2}}\;\mathrm{d}r\right)\;\mathrm{d}s,$